[Hallc_running] [New Logentry] Hall C RC Update (Sept 11)

brads at jlab.org brads at jlab.org
Fri Sep 11 15:45:01 EDT 2020

Logentry Text:
<p><a class="new" title="RC Update and Run plan for Sep 11,2020" href="https://hallcweb.jlab.org/wiki/index.php?title=RC_Update_and_Run_plan_for_Sep_11,2020">RC Update and Run plan for Sep 11,2020</a></p><h1><span id="RC_Meeting_Information" class="mw-headline">RC Meeting Information</span></h1><p><strong>RC Daily Meetings are M-F at 4:15 PM</strong></p><pre>To join the Meeting:  <a class="external free" href="https://bluejeans.com/861439920" rel="nofollow">https://bluejeans.com/861439920</a>
To join via phone :  +1.408.740.7256 (US (San Jose)), +1.888.240.2560 (US Toll Free))
</pre><h1><span id="General_updates" class="mw-headline">General updates</span></h1><p>(See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020">d2n Run Plan 2020</a> for broader status/overview.)</p><ul><li>Production running on Transverse target (@ 90); Kin-C, Kin-Z continues.<ul><li>General plan when this is done is to complete Kin-A, Kin-X (partially done already); then move on to Kin-B/Y and/or 1-pass running as time allows.</li></ul></li><li><strong>Moller</strong> measurement was rescheduled for today (2nd attempt); still working on it as of 3:30pm (Accelerator issues) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846082" rel="nofollow">[1]</a></li><li>Target was rotated to 90 deg yesterday</li><li>Max Target Tommy polarization @ 90 since yesterday has been 38.5% <a class="external autonumber" href="https://logbooks.jlab.org/entry/3845852" rel="nofollow">[2]</a><ul><li>For comparison, maximum seen with !
 target at 90 deg last week was 44% (9/5/2020 22:56:38)</li><li>Maximum seen at 270 was 37.5% <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844366" rel="nofollow">[3]</a>; Recent max around 36% (Thursday) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3845243" rel="nofollow">[4]</a></li><li>QWP optics were inspected during controlled access Friday morning and camera was re-aligned to get a better visual of the transverse QWP rotation stages <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846008" rel="nofollow">[5]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846026" rel="nofollow">[6]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846028" rel="nofollow">[7]</a></li></ul></li><li>Took ~8 hours of data in Swing, early OWL on Transverse @ 90 before machine fell over<ul><li>Lost beam around 2am in OWL with persistent vacuum fault (2L17), RF issue (2L18), !
 5th pass separator(s), etc <a class="external autonumber" href!
 ="https://logbooks.jlab.org/entry/3845800" rel="nofollow">[8]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3845845" rel="nofollow">[9]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846022" rel="nofollow">[10]</a></li><li>Trips not great during that period (35/hour), got a little better later in swing before beam went off completely in Owl</li></ul></li><li>Mike Nycz noticed that the SHMS s2Y(+/-) scalers stopped counting as of run 11247 (2020-09-03 21:00; last Thursday)<ul><li>Did some troubleshooting last night. ADC/TDC data seem OK, likely a bad module. <a class="external autonumber" href="https://logbooks.jlab.org/entry/3845681" rel="nofollow">[11]</a></li><li>Added scaler check to shift checklist <a class="external autonumber" href="https://logbooks.jlab.org/entry/3845688" rel="nofollow">[12]</a></li><li>Bad module replaced during Friday Day shift <a class="external autonumber" href="https://logbooks.jlab.org/ent!
 ry/3846102" rel="nofollow">[13]</a></li></ul></li><li>SHMS dipole tripped 2pm Friday (lead flow fault) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846117" rel="nofollow">[14]</a>. Reset cleanly <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846129" rel="nofollow">[15]</a></li></ul><ul><li>Hall C Techs helped us make some shim blocks for the target cell mounting jig 'V'-blocks Fri morning. This allows us to go ahead with mounting Butterball.</li><li>Retrieved the unfinished spectrum analyzer mounting parts from the JLab machine shop and worked with Hall C Techs to complete the work Fri morning<ul><li>Parts look to be complete now, dry-fitted everything we could</li><li>Junhao is locating the last few screws, dowels, etc that we need</li></ul></li><li><strong>Plan is to install this spectrum analyzer collimator at next opportunity</strong><ul><li>Needs followup with RadCon support [Brad]</li></ul></li></ul><h2><span id="Notes_f!
 or_the_Future" class="mw-headline">Notes for the Future</span></h2><ul>!
 <li>Beam Studies planned for <strong>Monday</strong>@ 10am to 6pm (including recovery), Sept 15<ul><li>RadCon will support going to CA at 10am, we can stay in CA for roughly 2 hours; OPS wants Hall C dump available after that</li><li>Simona wants access at this time for SOLID cherenkov related work</li><li>Target work? (install spectrum analyzer collimator if it can be ready)</li></ul></li></ul><ul><li><a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020#1-pass_Running">1-pass running</a> tentatively scheduled for Sept 17/18</li></ul><h2><span id="Pending_Issues" class="mw-headline">Pending Issues</span></h2><ul><li>Spectral analyzer collimator parts are complete (verify!)<ul><li>Junhao is locating the remaining screws, dowels, etc that will be needed for final mounting</li></ul></li></ul><ul><li>Target polarization for Tommy @ 270 is low (mid to high 30s)<ul><li>Spectrum analyzer collimator was broken at some point and present optic!
 s setup may be allowing transient light to interfere with spectrum measurement and QWP scans</li><li>JP wonders if the field direction is off for the transverse settings (Bill, Murchhana: see note in Target group chat from late Thursday)</li></ul></li></ul><ul><li>Beam energy drifts<ul><li>Hall C 5-pass polarization goes down as energy (HALLC:p) goes up <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844493" rel="nofollow">[16]</a></li></ul></li><li>Trip rates<ul><li>Hall A Limited to 140 uA to Reduce ILM0L02 Trips <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844844" rel="nofollow">[17]</a></li></ul></li></ul><ul><li>pNMR<ul><li>Quite different values than for NMR?</li><li>Summary for Tommy at 270 <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844699" rel="nofollow">[18]</a>, at 90 <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844044" rel="nofollow">[19]</a></li></ul></li></ul><h2>!
 <span id="Opportunistic_Jobs.2FAccess" class="mw-headline">Opportunisti!
 c Jobs/Access</span></h2><ul><li>Beam Studies planned for <strong>Monday</strong>, Sept 14 (Note date change!)<dl><dt>Target calibrations / tests</dt></dl><ul><li>Just before Moller measurement, and/or other known downtimes</li><li>See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020#Target_Calibrations_and_Tests">D2n_Run_Plan_2020#Target_Calibrations_and_Tests</a></li></ul></li></ul><ul><li><dl><dt>SOLID Cherenkov Test -- Access request [Simona]</dt></dl><ul><li>She would prefer to have access as soon as possible starting this Friday (Sep 11, Day Shift), to change the gas in the Cherenkov Tank during an opportunistic controlled access to the Hall during unexpected down times. NOTE: She may not need tech support, if she can use the ladder to reach fill the tank. \</li><li>Need ~1 hr without tech support. With tech support, may be 1-2 hrs.</li><li>Preferably, this work can be done during Day Shift, but it can also be done on Swing!
 /Owl.)</li></ul></li></ul><h1><span id="Current_Run_Plan" class="mw-headline">Current Run Plan</span></h1><h2><span id="Friday_DAY_.E2.86.92_Friday_SWING" class="mw-headline">Friday DAY → Friday SWING</span></h2><ul><li>(No beam since 2am...)</li><li><strong>Moller measurement</strong>planned for Friday DAY shift to begin around 9:30am<ul><li>Acclerator issues pushed this to 1pm</li></ul></li><li>Before handing over to Moller, the target experts will<ul><li>Take NMR/pNMR</li><li>Take escorted access to inspect QWP situation (ended up being opportunistic access)</li></ul></li><li>Wait for Moller to complete and get production beam re-established (late-DAY, early-SWING)</li></ul><ul><li><strong>Continue</strong> production data (HMS Kin-C/SHMS Kin-Z) with pol 3He target, <strong>Transverse/90</strong> polarization with NMR every 5 hrs.</li></ul><h2><span id="Friday_OWL_.E2.86.92_Thursday_OWL" class="mw-headline">Friday OWL → Thursday OWL</span></h2><ul><li>Continue produc!
 tion data (HMS Kin-C/SHMS Kin-Z) with pol 3He target, <strong>Transvers!
 e/90</strong> polarization with NMR every 5 hrs.</li></ul>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3846160
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200911/97a92f85/attachment-0001.html>

More information about the Hallc_running mailing list